You need to have 49,196 dollars to buy a new car 8 years from now. How much do you need to save at the end of each month if your savings account pays 8% per year, compounded quarterly? (note: round your answer to the nearest cent, and do not include spaces, currency signs, plus or minus signs, or commas)
Answer #1

Let the monthly saving be A

Quarterly savings = 3A

Quarterly rate = 8/4 = 2%

Future value = 49196

n = 8 years = 8*4 = 32 quarters

3A = 49196(A/F,2%,32)

3A = 49196*0.022611

3A = 1112.370756

A = 1112.370756/3

A = 370.79